OpenStudy (anonymous):

I really need help! I do not understand this. Question: choose one of the factors of 6x^3+6 a) x-1 b) x+1 c) x^2 -2x+1 d) x^2 + x+ 1 can you put the steps? I tried and I got x^3 + 3^3 = (x+3) (x^2 -3x+3^2) None of the two are in the answer choices so I know it's wrong! Help please.

OpenStudy (phi):

if you factor out the 6 you get 6(x^3 + 1) both x^3 and 1^3 are cubes (notice we can say 1= 1^3) see http://www.purplemath.com/modules/specfact2.htm for how to factor a sum of cubes

OpenStudy (anonymous):

I got as far as 6(x^3 + 1) but I get mixed up when plugging the terms in to a^3 +b^3 = (a+b) (a^2 -ab+6^2)

OpenStudy (anonymous):

its b) x=1

OpenStudy (phi):

to use this formula a^3 +b^3 match it with x^3 + 1^3 the a matches x and b matches 1 the formula says a^3 +b^3 = (a+b) (a^2 -ab+b^2) replace a with x, and b with 1 x^3 + 1^3 = (x+1)(x^2 -x*1 + 1^2) x*1 is just x, and 1^2 is 1*1 = 1 so we can simplify it to x^3 + 1^3 = (x+1)(x^2 -x + 1) I would say (x+1) is the answer.